OpenStudy (anonymous):

How do you solve cos(theta)+1=0

jimthompson5910 (jim_thompson5910):

cos(theta)+1=0 cos(theta)+1-1=0-1 cos(theta) = -1 Now use the unit circle to find all values of theta that make cos(theta) = -1 true

OpenStudy (anonymous):

Well im confused on the part of the unit circle and how to find the answer?

jimthompson5910 (jim_thompson5910):

do you have the unit circle with you

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

ok the x coordinate of any point represents the cosine of the angle for instance, at the point (1,0), the angle here is theta = 0 so this means cos(theta) = cos(0) = 1 sin(theta) = sin(0) = 0

jimthompson5910 (jim_thompson5910):

another point: at (0,1), theta is 90 degrees, which means cos(theta) = cos(90) = 0 sin(theta) = sin(90) = 1

jimthompson5910 (jim_thompson5910):

see how I'm getting these?

OpenStudy (anonymous):

So would the answer be 180

jimthompson5910 (jim_thompson5910):

good, and if you're restricting theta to be an angle between 0 and 360, then that's the only solution

jimthompson5910 (jim_thompson5910):

otherwise, it is theta = 180 + 360*n where n is any integer

OpenStudy (anonymous):

Okay thank you so im doing summer homework and would inwrite the answer like: cos(theta)+1=180

jimthompson5910 (jim_thompson5910):

no your answer needs to have theta isolated

jimthompson5910 (jim_thompson5910):

because that's exactly what it means to solve for a variable

OpenStudy (anonymous):

Ok so theta=180

jimthompson5910 (jim_thompson5910):

you're looking for values of theta that make cos(theta)+1=0 true if theta is between 0 and 360, then the only answer is theta = 180 if theta is not restricted, then the answers are theta = 180, theta = 540, theta = 900, etc etc

OpenStudy (anonymous):

Ok thank you! Also for tan^2(theta)-5=0 what wouldbyou do with the squared

jimthompson5910 (jim_thompson5910):

you would take the square root of both sides but this is after you add 5 to both sides

OpenStudy (anonymous):

So wouldmit look like this: theta= square root of 5

jimthompson5910 (jim_thompson5910):

if x^2 = 4, then x = 2 or x = -2 notice how both solutions work

OpenStudy (anonymous):

Yes sorry i accidentally closed the question

jimthompson5910 (jim_thompson5910):

when x = 2 x^2 = 2^2 = 4 and when x = -2, x^2 = (-2)^2 = 4 so that explains the plus/minus when dealing with square roots